OpenStudy (anonymous):

PLEASE HELP! Medals will be rewarded. Mary can paint a house 3 times faster than Amy can, working alone. If they can paint the house together in 9 hours, how long does it take Mary working alone?

OpenStudy (anonymous):

You can find similar problems to yours in http://www.saab.org/mathdrills/workcompletion.cgi

OpenStudy (anonymous):

Thanks. but it doesn't help

OpenStudy (anonymous):

\[m=3a\\ \frac{1}{a}+\frac{1}{m}=\frac{ 1}{9} \] These are the two equations you have to find. If you do a couple of problems on the link I gave you will understand how I came up with the second equation

OpenStudy (anonymous):

When you solve them, you get a=12 and m=36

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

I see how you got the fraction. But can you show me what to do afterwards step by step?

OpenStudy (anonymous):

Do you understand the second equation?

OpenStudy (anonymous):

yeah. the fractions are the amounts they can do in an hr

OpenStudy (anonymous):

\[m=3a\\ \frac{1}{a}+\frac{1}{m}=\frac{1}{9}\\ \frac{1}{a}+\frac{1}{3a}=\frac{1}{9}\\ \frac{3}{3a}+\frac{1}{3a}=\frac{1}{9}\\ \frac{4}{3a}=\frac{1}{9}\\ 36 =3a\\ a=12\\ m= 3a= 3(12)=36 \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

frac{3}{3a}+\frac{1}{3a}=\frac{1}{9} where did you get 3/3a?

OpenStudy (anonymous):

\[ \frac 1 a = \frac 3 {3a} \] Multiply up and down by 3